OpenStudy (anonymous):

You weight 300 N on Earth. A planet has three times the mass of Earth and four times the radius of the Earth. What is your weight in newtons on this planet

OpenStudy (anonymous):

Are you familiar with Newton's universal law of gravitation? \[F= \frac{GmM}{r^2}\] All you have to do is to do a comparison. Do you need further help?

OpenStudy (anonymous):

Sweet. I posted this awhile ago. Thanks Shadow

OpenStudy (anonymous):

You're welcome :) Your question is similar to his.

OpenStudy (anonymous):

Yes i'm familiar with it

OpenStudy (anonymous):

I did a ratio rather the gravitational force equation and solved for weight... but since I don't have mass for earth?

OpenStudy (anonymous):

Shadow, if possible, can you help me on my own question? I do not want to further de-rail Swin's question.

OpenStudy (anonymous):

\[ F_{earth} = \frac{Gm_{object} M_{earth}}{r^2_{earth}}\]---(1) \[ F_{planet} = \frac{Gm_{object} M_{planet}}{r^2_{planet}}\]---(2) Do a ratio of \(\frac{(1)}{(2)}\) and you should be able to get the answer.

OpenStudy (anonymous):

So I find two ratios?

OpenStudy (anonymous):

Yup. it's one actually. so \(\frac{F_{earth}}{F_{planet}}\)=?

OpenStudy (anonymous):

So I'm confused. They give you the weight on earth. 3X is the planet's mass (of the earth) and the radius is 4X the earth. But that equation solves for the gravitational force? ugh idk, maybe I'm overthinking

OpenStudy (anonymous):

Yup. weight on earth=\( F_{earth}\) where \(M_{planet}=3M_{earth} \)and so on

OpenStudy (anonymous):

The gravitational pull of earth is your weight. so g is a variable.

OpenStudy (anonymous):

Yeah, but what do I do with the 300 N?

OpenStudy (anonymous):

or is the mass for Earth can be constant?

OpenStudy (anonymous):

*any constant

OpenStudy (anonymous):

300N = \(F_{earth}\) M of earth , well, is a constant.

OpenStudy (anonymous):

300 N/ Fplanet (3*1/(4^2)??

OpenStudy (anonymous):

@Shadowys ? I feel like I'm doing this wrong lol

OpenStudy (anonymous):

did you mean \(\frac{300N}{F_{planet}}= \frac{4^2}{3}\)?

OpenStudy (anonymous):

Lol yess!!! I just put 300 N over a fraction/fraction :/ but yayyy right direction Thanks!!